6.
In a triangle, the sum of the interior angles is 180 degrees. If angle A is three times as large as angle B, and angles B and C sum to 90 degrees, what is the measure of each angle?

Answer by checkley77(12844)

A=3B GIVEN
A+B=90
REPLACE A WITH (3B) & SOLVE FOR B.
3B+B=90
4B=90
B=90/4
B=2.5 ANS.
A=3*22.5
A=67.5 ANS.
PROOF:
22.5+67.5=90
90=90